NevaehJade6219:

I need help with algebra! A model rocket is launched with an initial upward velocity of 235 ft/s. The rocket's height h (in feet) after t seconds is given by the following. h=235t-16t^2 Find all values of t for which the rocket's height is 151 feet.

NevaehJade6219:

@darkknight

hhanan:

Delete all of your comments this user needs help and STOP

hhanan:

@jhonyy9

NevaehJade6219:

@jimthompson5910

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan Delete all of your comments this user needs help and STOP \(\color{#0cbb34}{\text{End of Quote}}\)

jimthompson5910:

h=235t-16t^2 151=235t-16t^2 .... replace h with 151 16t^2-235t+151 = 0 ... get everything to one side You'll then use the quadratic formula \(\Large t = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\) to find the two solutions for t. Plug in a = 16, b = -235 and c = 151

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 h=235t-16t^2 151=235t-16t^2 .... replace h with 151 16t^2-235t+151 = 0 ... get everything to one side You'll then use the quadratic formula \(\Large t = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\) to find the two solutions for t. Plug in a = 16, b = -235 and c = 151 \(\color{#0cbb34}{\text{End of Quote}}\) i need what t=

jimthompson5910:

If a = 16, b = -235 and c = 151 then what is \(b^2 - 4ac\) equal to?

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 If a = 16, b = -235 and c = 151 then what is \(b^2 - 4ac\) equal to? \(\color{#0cbb34}{\text{End of Quote}}\) IDK thats y im asking uuu

jimthompson5910:

we would replace the variables with the numbers they're equal to b^2 - 4ac = (-235)^2 - 4(16)(151) = ??

NevaehJade6219:

ok im gonna try it now

NevaehJade6219:

UGH IDK

jimthompson5910:

What kind of calculator do you have?

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 What kind of calculator do you have? \(\color{#0cbb34}{\text{End of Quote}}\) ti 84

jimthompson5910:

If you dont have one, you can use this free calculator https://www.desmos.com/calculator

jimthompson5910:

ok ti 84 works too

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 If you dont have one, you can use this free calculator https://www.desmos.com/calculator \(\color{#0cbb34}{\text{End of Quote}}\) lol i have that to

jimthompson5910:

whichever you prefer

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 whichever you prefer \(\color{#0cbb34}{\text{End of Quote}}\) ok im getting my ti

jimthompson5910:

alright

NevaehJade6219:

i have it

jimthompson5910:

ok type in ` (-235)^2 - 4(16)(151) `

NevaehJade6219:

i got 45561

jimthompson5910:

same here, now apply the square root to that \(\sqrt{45561} = \ ?\)

NevaehJade6219:

? wat

NevaehJade6219:

oohhhh i got 213.45

jimthompson5910:

correct

NevaehJade6219:

is that my answer?

jimthompson5910:

not yet

jimthompson5910:

So we have this so far \(\Large t = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\) \(\Large t = \frac{-(-235)\pm\sqrt{ (-235)^2 - 4(16)(151) }}{2(16)}\) \(\Large t = \frac{235\pm\sqrt{45561}}{32}\) \(\Large t \approx \frac{235\pm213.45}{32}\) \(\Large t \approx \frac{235+213.45}{32} \ \text{ or } t \approx \frac{235-213.45}{32}\) How can we finish up?

NevaehJade6219:

t=241.67 and t=228.33

jimthompson5910:

Not sure how you got those

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 Not sure how you got those \(\color{#0cbb34}{\text{End of Quote}}\) ugh i did 235+213.45/32

jimthompson5910:

oh you have to add first, then divide later

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 oh you have to add first, then divide later \(\color{#0cbb34}{\text{End of Quote}}\) so what do i add?

jimthompson5910:

The stuff up top is effectively in its own parenthesis grouping Meaning that \(\frac{235+213.45}{32}\) is the same as \((235+213.45)/32\) and similar for the minus portion as well So, \(\Large t \approx \frac{235+213.45}{32} \approx \frac{448.45}{32} \approx \ ?\)

NevaehJade6219:

t=14.01 t=5226

jimthompson5910:

I agree with t = 14.01 Not sure how you got the other value though

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @NevaehJade6219 t=14.01 t=5226 \(\color{#0cbb34}{\text{End of Quote}}\) **15.26 i mistyped

jimthompson5910:

What did you type in to get 15.26?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan @jhonyy9 \(\color{#0cbb34}{\text{End of Quote}}\) ty you tagged me but i ve answered this problem past 1-2 hours ago

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 What did you type in to get 15.26? \(\color{#0cbb34}{\text{End of Quote}}\) 488.45/32

jimthompson5910:

how did you get 488.45?

NevaehJade6219:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 how did you get 488.45? \(\color{#0cbb34}{\text{End of Quote}}\) ugh mistype again i need to redu that part it was really 448.45

NevaehJade6219:

so t=14.01and 14.01

jimthompson5910:

The other piece you need to evaluate is \(\large t \approx \frac{235-213.45}{32}\) You need to evaluate the stuff up top first, then divide later.

NevaehJade6219:

.67?

jimthompson5910:

Correct, so the two approximate solutions are t = 14.01 and t = 0.67 This means that the rocket reaches the height of 151 ft at around 0.67 seconds and around 14.01 seconds In this diagram  Point A represents when t = 0.67 Point B is when t = 14.01

NevaehJade6219:

so t=0.67 and t=14.01 is the answer?

jimthompson5910:

Yes there are two time values when the rocket is at this height

NevaehJade6219:

oki! thanks so much old man!